OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola -1/16x^2 = y

OpenStudy (mertsj):

\[-\frac{1}{16}(x-h)^2=y-k\]

OpenStudy (mertsj):

What is the vertex?

OpenStudy (anonymous):

0,0

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Multiply both sides by 16 and you will have it in the form:

OpenStudy (mertsj):

(x−h)^2=4p(y−k)

OpenStudy (anonymous):

Idk

OpenStudy (mertsj):

When you have identified 4p, you can find p. Remember p is the distance from the vertex to the focus and also the distance from the vertex to the directrix.

OpenStudy (mertsj):

You don't know what? How to multiply both sides by 16?

OpenStudy (anonymous):

What are the two sides?

OpenStudy (mertsj):

x=3 The two sides are: left side = s right side = 3 If I multiply both sides by 6, I will have 6x=18

OpenStudy (anonymous):

I thought by 16?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Did you multiply both sides by 16?

OpenStudy (anonymous):

16s = 48

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

Look at me.

OpenStudy (anonymous):

I am. I'm really trying :(

OpenStudy (mertsj):

Take your original equation that is given in the problem and multiply both sides by 16

OpenStudy (anonymous):

16y = -x^2

OpenStudy (mertsj):

now if you switch sides, it is of the form:

OpenStudy (mertsj):

\[-x^2=4py\]

OpenStudy (mertsj):

What is 4p?

OpenStudy (anonymous):

P = 4

OpenStudy (mertsj):

yes. p = 4. Does this one open up or down?